OpenStudy (ccastudentgrade10):

fan&medal :) simplify the expression -6+i/-5+i

OpenStudy (ccastudentgrade10):

@Vocaloid ?

OpenStudy (lochana):

you want to multiply denominator and numerator by -5 - i . so that denominator will no longer be complex number.

OpenStudy (ccastudentgrade10):

so 30 + i/25?

OpenStudy (lochana):

\[\frac{ (-6+i)(-5 - i) }{ (-5+i)(-5-i) }\] \[\frac{ 30 + 6i - 5i - i^2 }{ 25 - i^2 }\]

OpenStudy (lochana):

\[\frac{ 31 + i }{ 26 }\]

OpenStudy (lochana):

\[\frac{ 31}{ 26 }+ \frac{i}{26}\]

OpenStudy (ccastudentgrade10):

oh! i get it. thank you!!!

OpenStudy (lochana):

bye